OpenStudy (anonymous):

If someone swam 4 miles in 6 minutes, how many minutes wil it take for them to reach 10 miles?

OpenStudy (anonymous):

14

OpenStudy (texaschic101):

proportion time.. 4/6 = 10/x (4 miles to 6 min = 10 miles to x minutes) cross multiply (4)(x) = (10)(6) 4x = 60 x = 60/4 x = 15 minutes